# AeroSketch: Near-Optimal Time Matrix Sketch Framework for Persistent, Sliding Window, and Distributed Streams

### v1

@@ 4.3. Distributed Matrix Sketch (Problem 5)
1
2Procedure Update($\bm{a}_{i}$): // at Site $j$
3 $F_{j}\leftarrow F_{j}+\left\|\bm{a}_{i}\right\|_{2}^{2}$
4 if $F_{j}\geq\frac{\varepsilon}{m}\hat{F}$ then
5      Send $F_{j}$ to coordinator; set $F_{j}\leftarrow 0$
6
7 AeroSketch.Update($\bm{a}_{i}$)
@@ 5      Send $F_{j}$ to coordinator; set $F_{j}\leftarrow 0$
9      Send $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$ of $\mathcal{S}[0]$ to coordinator
@@
11    return $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$
12 case received $\hat{F}$ do
@@
14
15
16Procedure Receive_Update: // at Coordinator
17 case received $F_{j}$ do
18    $\hat{F}\mathrel{+}=F_{j}$ and $\texttt{msg}\mathrel{+}=1$
19    if $\texttt{msg}\geq m$ then
20       $\texttt{msg}\leftarrow 0$, then Broadcast $\hat{F}$ to all sites
21
22 case received $\bm{Z}$ and $\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$ do
23    $\bm{B}\mathrel{+}=\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}+\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}-\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}$
24
25
26
27Procedure Query: // at Coordinator
28 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
29 $\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$
30 return $\hat{\bm{B}}[1:\ell,:]$
31
@@ 28 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
We defer the proof of Theorem 4.3 to Appendix G.
@@ 4.3.1. Tracking Distributed Matrix Sketch over Sliding Window (Problem 6)
Algorithm 11 describes how P5 can be used to solve the Tracking Distributed Matrix Sketch over Sliding Window problem. The high-level idea and algorithm skeleton are similar to DA2 (Algorithm 5 in (Zhang et al., 2017)). The IWMT protocol used by DA2 is a “black-box” one-way communication version of P2. Here we replace IWMT with P5 to achieve the same communication complexity while avoiding frequent and time-consuming matrix factorizations.
1
2Procedure Update($\bm{a}_{i}$): // at Site $j$
3 $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\leftarrowP$5.Update($\bm{a}_{i}$)
4   Send $(\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C},\texttt{flag}=\texttt{update})$
5 $mEH_{a}.\textsc{Update}(\bm{a}_{i})$;
6 if $i\mod N\equiv 0$ then
7    $Q\leftarrow IWMT_{c}(mEH_{a})$
8 for $\bm{e}_{t}\in Q$ and t<iNt<i-N do
9      Send $(\bm{e}_{t},\texttt{flag}=\texttt{expire})$, then remove $\bm{e}_{t}$ from $Q$
10
11
12Procedure Receive_Update: // at Coordinator
13 switch flag do
14    case update do
15       $\bm{B}\mathrel{+}=\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}+\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}-\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}$
16
17    case expire do
18       $\bm{B}\mathrel{-}=\bm{e}_{t}^{\top}\bm{e}_{t}$
19
20
21
22Procedure Query: // at Coordinator
23 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
24 $\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$
25 return $\hat{\bm{B}}[1:\ell,:]$
@@ 23 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
We defer the proof of Theorem 4.4 to Appendix H.
@@ 5.2. Experimental Results
Figures 8 and 9 show the space/communication costs vs. error. We observe that, under the same empirical error, the space/communication costs of baseline algorithms and our AeroSketch are comparable. In some cases, AeroSketch even achieves lower space costs, consistent with our earlier theoretical analysis that its space complexity is of the same order and asymptotically optimal compared to baseline algorithms.
@@ Appendix I Optimize aDS-COD by our AeroSketchCOD Framework
2
3Initialize: Initialize a AeroSketchCOD (Algorithm 9) sketch $AS$ and an auxiliary sketch $AS_{\text{aux}}$ with initial threshold $\varepsilon N$ and threshold level $L=1$.
4
5Function Update($\bm{x}_{i},\bm{y}_{i}$):
6 while $AS.S[0].t+N\leq i$ do
7    $AS.S.\operatorname{Dequeue}()$;
8   end while
9
10 if $i\mod N==1$ then
11    $AS\leftarrow AS_{\text{aux}}$;
12    $AS_{\text{aux}}.\operatorname{Initialize}()$;
13   end if
14
15 $AS.\operatorname{Update}(\bm{x}_{i},\bm{y}_{i})$;
16 $AS_{\text{aux}}.\operatorname{Update}(\bm{x}_{i},\bm{y}_{i})$;
17 if $\operatorname{len}(AS.S)\geq\frac{L}{\varepsilon}$ then
18    $L\leftarrow L+1$;
19    $AS.\theta\leftarrow 2\cdot AS.\theta$;
20   end if
21 else if $\operatorname{len}(AS.S)\leq\frac{L-1}{\varepsilon}$ then
22    $L\leftarrow L-1$;
23    $AS.\theta\leftarrow AS.\theta/2$;
24   end if
25
26
27Function Query():
28 return $AS.\operatorname{Query}()$;
29

### v2

@@ 4.3. Distributed Matrix Sketch (Problem 5)
Procedure Update($\bm{a}_{i}$): // at Site $j$
 $F_{j}\leftarrow F_{j}+\left\|\bm{a}_{i}\right\|_{2}^{2}$
 if $F_{j}\geq\frac{\varepsilon}{m}\hat{F}$ then
      Send $F_{j}$ to coordinator; set $F_{j}\leftarrow 0$
   end if
 AeroSketch.Update($\bm{a}_{i}$)
@@
      Send $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$ of $\mathcal{S}[0]$ to coordinator
@@
    return $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$
   end if
 case received $\hat{F}$ do
@@
Procedure Receive_Update: // at Coordinator
 case received $F_{j}$ do
    $\hat{F}\mathrel{+}=F_{j}$ and $\texttt{msg}\mathrel{+}=1$
    if $\texttt{msg}\geq m$ then
       $\texttt{msg}\leftarrow 0$, then Broadcast $\hat{F}$ to all sites
      end if
 case received $\bm{Z}$ and $\bm{Z}^{\top}\bm{C}^{\top}\bm{C}$ do
    $\bm{B}\mathrel{+}=\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}+\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}-\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}$
Procedure Query: // at Coordinator
 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
 $\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$
 return $\hat{\bm{B}}[1:\ell,:]$
@@
The detailed proof can be found in Appendix G.
@@ 4.3.1. Tracking Distributed Matrix Sketch over Sliding Window (Problem 6)
Algorithm 11 describes how P5 (Algorithm 10) can be used to solve the Tracking Distributed Matrix Sketch over Sliding Window problem. The high-level idea and algorithm skeleton are similar to DA2 (Algorithm 5 in (Zhang et al., 2017)). The IWMT protocol used by DA2 is a “black-box” one-way communication version of P2. Here we replace IWMT with P5 (Algorithm 10) to achieve the same communication complexity while avoiding frequent and time-consuming matrix factorizations.
Procedure Update($\bm{a}_{i}$): // at Site $j$
 $\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\leftarrowP$5.Update($\bm{a}_{i}$)
   Send $(\bm{Z},\bm{Z}^{\top}\bm{C}^{\top}\bm{C},\texttt{flag}=\texttt{update})$
 $mEH_{a}.\textsc{Update}(\bm{a}_{i})$;
 if $i\mod N\equiv 0$ then
    $Q\leftarrow IWMT_{c}(mEH_{a})$
   end if
 for $\bm{e}_{t}\in Q$ and t<iNt<i-N do
      Send $(\bm{e}_{t},\texttt{flag}=\texttt{expire})$, then remove $\bm{e}_{t}$ from $Q$
   end for
Procedure Receive_Update: // at Coordinator
 switch flag do
    case update do
       $\bm{B}\mathrel{+}=\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}+\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}-\bm{Z}\bm{Z}^{\top}\bm{C}^{\top}\bm{C}\bm{Z}\bm{Z}^{\top}$
    case expire do
       $\bm{B}\mathrel{-}=\bm{e}_{t}^{\top}\bm{e}_{t}$
   end switch
Procedure Query: // at Coordinator
 $\bm{V},\bm{\Lambda},\bm{V}^{\top}]\leftarrow\mathtt{eigen\_decomposition}(\bm{B})$
 $\hat{\bm{B}}\leftarrow\sqrt{\max\left(\bm{\Lambda}-\bm{I}\lambda_{\ell},\bm{0}\right)}\,\bm{V}^{\top}$
 return $\hat{\bm{B}}[1:\ell,:]$
@@
The detailed proof can be found in Appendix H.
@@ 5.2. Experimental Results
Figures 8 and 9 show the space/communication costs vs. error. We observe that, under the same empirical error, the space/communication costs of baseline algorithms and our AeroSketch are comparable. In some cases, AeroSketch even achieves lower space costs, consistent with our earlier theoretical analysis that its space/communication complexity is of the same order and asymptotically optimal compared to baseline algorithms.
@@ Appendix I Optimize aDS-COD by our AeroSketchCOD Framework
Initialize: Initialize a AeroSketchCOD (Algorithm 9) sketch $AS$ and an auxiliary sketch $AS_{\text{aux}}$ with initial threshold $\varepsilon N$ and threshold level $L=1$.
Function Update($\bm{x}_{i},\bm{y}_{i}$):
 while $AS.S[0].t+N\leq i$ do
    $AS.S.\operatorname{Dequeue}()$;
   end while
 if $i\mod N==1$ then
    $AS\leftarrow AS_{\text{aux}}$;
    $AS_{\text{aux}}.\operatorname{Initialize}()$;
   end if
 $AS.\operatorname{Update}(\bm{x}_{i},\bm{y}_{i})$;
 $AS_{\text{aux}}.\operatorname{Update}(\bm{x}_{i},\bm{y}_{i})$;
 if $\operatorname{len}(AS.S)\geq\frac{L}{\varepsilon}$ then
    $L\leftarrow L+1$;
    $AS.\theta\leftarrow 2\cdot AS.\theta$;
   end if
 else if $\operatorname{len}(AS.S)\leq\frac{L-1}{\varepsilon}$ then
    $L\leftarrow L-1$;
    $AS.\theta\leftarrow AS.\theta/2$;
   end if
Function Query():
 return $AS.\operatorname{Query}()$;